In a given population of flies, pink skin (e) is recessive to the wild-type skin (E). A population has 96 pink flies, and 125 that have wild type skin. What is the frequency of alleles in this population? Assume HW equilibrium.

Explain every step please
Answer #1

total number of flies=96+125

=221

Frequency of pink flies ( genotypic frequency of ee)= number of pink flies/ total number

=96/125

=0.768

frequency of allele e= square root of the genotypic frequency of ee

= square root of 0.768

=0.8763

frequency of allele E+ frequency of allele e=1

frequency of allele E= 1- frequency of allele e

=1-0.8763

=0.1237

.